# Distributed Model-Predictive Energy Management Strategy for Shipboard Power Systems Considering Battery Degradation

### v1

@@
Abstract
@@
keywords:
Battery degradation, model predictive control, energy management
thanks: This work is sponsored by U.S Office of Naval Research grants N00014-22-1-2442, N00014-16-1-2956 and N00014-21-1-2124.thanks: DISTRIBUTION STATEMENT A. APPROVED for public release, distribution is unlimited. Approved, DCN# 0543-2005-24.
@@ 1 Introduction
Conceptualization of Microgrids (MGs) as an amalgamation of fossil-fuel-driven micro-turbines, renewable energy-based generating systems, and energy storage systems (ESSs) incorporated into a distribution generation environment, driving high demand for the power. MGs are categorized into grid-connected and islanded, based on the configuration and their mode of operation (Hatziargyriou et al. (2007)). One such example of an islanded operation is Shipboard Power Systems (SPSs). In shipboard power systems, power generation modules (PGMs) and power conversion modules (PCMs) provide the power required by the loads through means of the direct current (DC) distribution system. Modern SPSs are equipped with advanced power loads such as rail guns, electromagnetic radars, and heavy nonlinear and pulsed power loads (PPLs) (Doerry and Amy (2015)).
There is a challenge to utilize the existing PGMs, which are ramp-rate limited, to meet the power demanded by the high-ramp-rate loads. Failure to ramp up in appropriate time to meet the load requirements can lead to system imbalances and instability. Furthermore, adding additional generators is not a plausible solution considering the ship hull dimension restrictions. This imposes a need for high-ramp-rate power generation elements to be integrated into SPSs. Energy storage systems such as batteries can support higher ramp rates. Integrating energy storage systems (ESSs) into the existing notional SPS framework addresses the high ramp requirements (Team (2020)). The ESSs can store the energy during normal periods and when needed can provide immediate ramp-up support. Several investigators have addressed the above challenge without considering battery degradation (Vu et al. (2017); Bijaieh et al. (2021); Seenumani et al. (2008)) where they have used dynamic programming (DP) and model predictive control (MPC) to address the aforementioned challenge. MPC is a renowned concept used in control literature whose applications include industry. MPC problem solves an objective function subject to a set of constraints that can be actuator limitations in physical systems or rate constraints (Wang and Boyd (2010)). The main advantage of the MPC over DP is that it enables incorporation of the constraints such as inequality and box constraints capturing the upper and lower limitations and ramp rates, thus accounting for actuator limitations.
@@
Nevertheless, batteries degrade faster than generators. This fact imposes another challenge to consider both ramp rate requirements and battery degradation together. Some investigators have considered this challenge while addressing the EM problem for the MGs (Hein et al. (2021); Antoniadou-Plytaria et al. (2021); Li et al. (2023); Zhao and Li (2024); Nawaz et al. (2023); Ji et al. (2023); Wang et al. (2022); Vedula et al. (2021)). In Hein et al. (2021), the authors use DP to solve the EM problem while considering the depth of discharge (DoD) as a measure for capturing battery degradation. In Antoniadou-Plytaria et al. (2021), an energy market pricing-based EM problem for an MG, considering the battery DoD as a degradation measure, is presented. Market-energy-pricing-based approach for MG with fast Li-ion(standing for Lithium-ion)-based batteries, considering DoD as a measure for capturing the battery degradation, has been presented in Li et al. (2023). DoD has been also used as a battery degradation measure to solve the EM problem in Zhao and Li (2024) where the authors proposed a neural-network-based method to address the EM problem in the MGs. In Nawaz et al. (2023), the authors have presented a state-of-charge (SoC)-based EM scheme for MGs; the SoC is regulated around a reference value to minimize battery degradation. Monitoring the battery charge/discharge cycles and using them to determine the battery degradation with an application to the MGs has been investigated in Ji et al. (2023). The work of Hein et al. (2021); Antoniadou-Plytaria et al. (2021); Li et al. (2023); Zhao and Li (2024); Nawaz et al. (2023); Ji et al. (2023) consider battery aging in different measures such as DoD, SoC, and charge/discharge based cycle count. In Wang et al. (2022), the measures used to minimize the battery usage are battery power and battery SoC, but the authors in Wang et al. (2022) did not consider the cost associated with the generators and its effect on determining the battery degradation.
@@ 3 Model Development
Since the design and analysis of such a DLC is not the main objective of this paper, numerous linear (PID) and nonlinear methods can be employed to design it (Khalil (2014)).
@@ 4 Proposed Energy Management Method
where $k\triangleq\left[\begin{array}[]{cccc}1&2&\ldots&h\end{array}\right]^{\top}\in\mathbb{R}^{h}$ is the length of the prediction horizon, $\mathbf{p}_{g_{i}}\triangleq\left[\begin{array}[]{cccc}{\mathbf{p}_{g_{i}}}_{k}&{\mathbf{p}_{g_{i}}}_{k+1}&\ldots&{\mathbf{p}_{g_{i}}}_{k+h-1}\end{array}\right]^{\top}\in\mathbb{R}^{h}$ is the power profile for generator $i$ over the prediction horizon of length $h$, $\textbf{p}_{b_{j}}\triangleq\left[\begin{array}[]{cccc}{\mathbf{p}_{b_{j}}}_{k}&{\mathbf{p}_{b_{j}}}_{k+1}&\ldots&{\mathbf{p}_{b_{j}}}_{k+h-1}\end{array}\right]^{\top}\in\mathbb{R}^{h}$ is the power profile for battery $j$ over the prediction horizon of length $h$ and ${p}_{f}\mathbf{1}\in\mathbb{R}^{h}$ is the desired total power held fixed over the prediction horizon. The generator cost ${C}_{g_{i}}:\mathbb{R}^{{h}}\longrightarrow\mathbb{R}_{+}$ could be associated with an efficiency map or operating at a desired rated power. The battery cost ${C}_{b_{j}}:\mathbb{R}^{{h}}\longrightarrow\mathbb{R}_{+}$ could be the cost associated with battery health monitoring and degradation management. ${h}\in\mathbb{N}$ represents the prediction horizon. $\mathcal{X}_{g}\subset\mathbb{R}^{h}$ and $\mathcal{X}_{b}\subset\mathbb{R}^{h}$ represents the set of inclusion constraints each of the PGM and PCM summarizing attributes such as upper and lower power limits, ramp rate limits, and relevant system dynamics. The optimization variables are optimized for the entire length of the horizon and the first value of the sequence is supplied/applied as a power reference (control input) to the low-level (DLC). The PGM and the PCM costs are considered so that the objective is to maintain the PGM around a rated value ($\mathbf{p}_{g}^{r}$) which is known and to minimize the PCM power. Based on the optimization problem presented in (7), consider the following MPC problem with the respective cost functions for the PGM and the PCM as follows:
@@ 4.1 Distributed Formulation
where $\boldsymbol{\lambda}\in\mathbb{R}^{h}$ is the dual variable associated with the main problem in (7) and $\mathcal{I_{X}}_{g}$ and $\mathcal{I_{X}}_{b}$ are the indicator functions for the inclusion constraints. Given $\boldsymbol{\lambda}$, let $\mathbf{p}_{g_{i}}^{*}$ and $\mathbf{p}_{b_{j}}^{*}$ to be the solutions to the optimization problem in (7) (Boyd and Vandenberghe (2004))
@@
for a step size $\alpha>0$ and $t$ is the iteration counter. Since the minimizing power profiles $\textbf{p}^{*}_{g_{i}}$ and $\textbf{p}^{*}_{b_{j}}$ depends on $\boldsymbol{\lambda}$, the dual ascent step in (10) becomes a recursion and challenging to evaluate. Consequently, we incorporate the primal problems in (11) within a sequential minimization-maximization scheme Boyd et al. (2011) as follows:
@@ 4.3 PCM Node Optimization problem
where, $\kappa={T_{d}}/{Qv_{bus}}$, $r_{b}\in\mathbb{R}_{+}$ is the ramp-rate limitation associated with battery, $\overline{\mathbf{p}}_{b_{j}}\in\mathbb{R}_{+}$ and $\underline{\mathbf{p}}_{b_{j}}\in\mathbb{R}_{-}$ are the upper and lower limits on battery power. This indicates the discharging and the charging mode of the PCM. The solution to the optimization problems in (12) and (13) are used in calculating the dual variable update. The schematic for this aggregation and distribution setup is provided in Fig. 3. All the nodes present in the system send their current iterate value to the dual updater node and the information of the previous iterate of the dual variable $\boldsymbol{\lambda}_{t}$ is sent to all the nodes. Thus, the dual updater, which acts as a common scheduler to all nodes implements the update law given in (11c). The convergence of the proposed algorithm under the assumed objectives and the constraint sets which are polytopes is well established in the literature (Boyd and Vandenberghe (2004); Boyd et al. (2011)).
@@ 5 Numerical Simulation
The designed model predictive energy management strategy is tested on a single PGM, PCM, and Load scenario to fully study and analyze the effect of the weights $\beta$ and $\gamma$ on the total energy delivered by the PCM, PGM, and its effect on the PCM capacity loss %. The simulation environment used in this setup is based on MATLAB-Simulink on a real-time simulator. The simulation is run on a desktop with the following configurations Digital Storm Intel Core i9-13900K (5.7GHz Turbo) with 64GB RAM acting as a Host PC and a real-time target with Intel Core 3.6GHz 8-core and 32GB RAM (shown in Fig. 3). The optimization problem presented in (8) with $n_{g}=1$ and $n_{b}=1$ is implemented to study the trade-off characteristics. The parameters used are based on the documentation and the notional SPS component sizing provided by the Office of Naval Research (ONRs) ESRDC (Team (2020)). For all the results presented, the prediction horizon is fixed at ($h=5$seconds). The simulation was run at the fixed step of $10^{-3}$seconds. The rate transition or the communication (measurement exchange) delay between the model and the MPC EM layer was set at $1$second to facilitate optimization convergence.

### v2

@@
Distributed Model-Predictive Energy Management Strategy for Shipboard Power Systems Considering Battery Degradation
Satish Vedula, Seyyed Shaho Alaviani∗∗ and
Olugbenga Moses Anubi
* Department of Electrical and Computer Engineering, the Center for Advanced Power Systems, Florida State University E-mail: {svedula , oanubi}@fsu.edu **∗ ∗ Center for Advanced Power Systems (CAPS), Florida State University, and National High Magnetic Field Laboratory (MagLab), Tallahassee, FL 32310, USA (email: salaviani@fsu.edu)
1 Abstract
@@
Keywords: Battery degradation, model predictive control, energy management
@@ 2 Introduction
Conceptualization of Microgrids (MGs) as an amalgamation of fossil-fuel-driven micro-turbines, renewable energy-based generating systems, and energy storage systems (ESSs) incorporated into a distribution generation environment, driving high demand for the power. MGs are categorized into grid-connected and islanded, based on the configuration and their mode of operation ([1]). One such example of an islanded operation is Shipboard Power Systems (SPSs). In shipboard power systems, power generation modules (PGMs) and power conversion modules (PCMs) provide the power required by the loads through means of the direct current (DC) distribution system. Modern SPSs are equipped with advanced power loads such as rail guns, electromagnetic radars, and heavy nonlinear and pulsed power loads (PPLs) ([2]).
There is a challenge to utilize the existing PGMs, which are ramp-rate limited, to meet the power demanded by the high-ramp-rate loads. Failure to ramp up in appropriate time to meet the load requirements can lead to system imbalances and instability. Furthermore, adding additional generators is not a plausible solution considering the ship hull dimension restrictions. This imposes a need for high-ramp-rate power generation elements to be integrated into SPSs. Energy storage systems such as batteries can support higher ramp rates. Integrating energy storage systems (ESSs) into the existing notional SPS framework addresses the high ramp requirements ([3]). The ESSs can store the energy during normal periods and when needed can provide immediate ramp-up support. Several investigators have addressed the above challenge without considering battery degradation ([4, 5, 6]) where they have used dynamic programming (DP) and model predictive control (MPC) to address the aforementioned challenge. MPC is a renowned concept used in control literature whose applications include industry. MPC problem solves an objective function subject to a set of constraints that can be actuator limitations in physical systems or rate constraints ([7]). The main advantage of the MPC over DP is that it enables incorporation of the constraints such as inequality and box constraints capturing the upper and lower limitations and ramp rates, thus accounting for actuator limitations.
@@
Nevertheless, batteries degrade faster than generators. This fact imposes another challenge to consider both ramp rate requirements and battery degradation together. Some investigators have considered this challenge while addressing the EM problem for the MGs ([8, 9, 10, 11, 12, 13, 14, 15]). In [8], the authors use DP to solve the EM problem while considering the depth of discharge (DoD) as a measure for capturing battery degradation. In [9], an energy market pricing-based EM problem for an MG, considering the battery DoD as a degradation measure, is presented. Market-energy-pricing-based approach for MG with fast Li-ion(standing for Lithium-ion)-based batteries, considering DoD as a measure for capturing the battery degradation, has been presented in [10]. DoD has been also used as a battery degradation measure to solve the EM problem in [11] where the authors proposed a neural-network-based method to address the EM problem in the MGs. In [12], the authors have presented a state-of-charge (SoC)-based EM scheme for MGs; the SoC is regulated around a reference value to minimize battery degradation. Monitoring the battery charge/discharge cycles and using them to determine the battery degradation with an application to the MGs has been investigated in [13]. The work of [8, 9, 10, 11, 12, 13] consider battery aging in different measures such as DoD, SoC, and charge/discharge based cycle count. In [14], the measures used to minimize the battery usage are battery power and battery SoC, but the authors in [14] did not consider the cost associated with the generators and its effect on determining the battery degradation.
@@ 4 Model Development
Since the design and analysis of such a DLC is not the main objective of this paper, numerous linear (PID) and nonlinear methods can be employed to design it ([16]).
@@ 5 Proposed Energy Management Method
where $k\triangleq\left[\begin{array}[]{cccc}1&2&\ldots&h\end{array}\right]^{\top}\in\mathbb{R}^{h}$ is the length of the prediction horizon, $\mathbf{p}_{g_{i}}\triangleq\left[\begin{array}[]{cccc}{\mathbf{p}_{g_{i}}}_{k}&{\mathbf{p}_{g_{i}}}_{k+1}&\ldots&{\mathbf{p}_{g_{i}}}_{k+h-1}\end{array}\right]^{\top}\in\mathbb{R}^{h}$ is the power profile for generator $i$ over the prediction horizon of length $h$,
$\textbf{p}_{b_{j}}\triangleq\left[\begin{array}[]{cccc}{\mathbf{p}_{b_{j}}}_{k}&{\mathbf{p}_{b_{j}}}_{k+1}&\ldots&{\mathbf{p}_{b_{j}}}_{k+h-1}\end{array}\right]^{\top}\in\mathbb{R}^{h}$ is the power profile for battery $j$ over the prediction horizon of length $h$ and ${p}_{f}\mathbf{1}\in\mathbb{R}^{h}$ is the desired total power held fixed over the prediction horizon. The generator cost ${C}_{g_{i}}:\mathbb{R}^{{h}}\longrightarrow\mathbb{R}_{+}$ could be associated with an efficiency map or operating at a desired rated power. The battery cost ${C}_{b_{j}}:\mathbb{R}^{{h}}\longrightarrow\mathbb{R}_{+}$ could be the cost associated with battery health monitoring and degradation management. ${h}\in\mathbb{N}$ represents the prediction horizon. $\mathcal{X}_{g}\subset\mathbb{R}^{h}$ and $\mathcal{X}_{b}\subset\mathbb{R}^{h}$ represents the set of inclusion constraints each of the PGM and PCM summarizing attributes such as upper and lower power limits, ramp rate limits, and relevant system dynamics. The optimization variables are optimized for the entire length of the horizon and the first value of the sequence is supplied/applied as a power reference (control input) to the low-level (DLC). The PGM and the PCM costs are considered so that the objective is to maintain the PGM around a rated value ($\mathbf{p}_{g}^{r}$) which is known and to minimize the PCM power. Based on the optimization problem presented in (7), consider the following MPC problem with the respective cost functions for the PGM and the PCM as follows:
@@ 5.1 Distributed Formulation
where $\boldsymbol{\lambda}\in\mathbb{R}^{h}$ is the dual variable associated with the main problem in (7) and $\mathcal{I_{X}}_{g}$ and $\mathcal{I_{X}}_{b}$ are the indicator functions for the inclusion constraints. Given $\boldsymbol{\lambda}$, let $\mathbf{p}_{g_{i}}^{*}$ and $\mathbf{p}_{b_{j}}^{*}$ to be the solutions to the optimization problem in (7) ([18])
@@
for a step size $\alpha>0$ and $t$ is the iteration counter. Since the minimizing power profiles $\textbf{p}^{*}_{g_{i}}$ and $\textbf{p}^{*}_{b_{j}}$ depends on $\boldsymbol{\lambda}$, the dual ascent step in (10) becomes a recursion and challenging to evaluate. Consequently, we incorporate the primal problems in (11) within a sequential minimization-maximization scheme [19] as follows:
@@ 5.3 PCM Node Optimization problem
where, $\kappa={T_{d}}/{Qv_{bus}}$, $r_{b}\in\mathbb{R}_{+}$ is the ramp-rate limitation associated with battery, $\overline{\mathbf{p}}_{b_{j}}\in\mathbb{R}_{+}$ and $\underline{\mathbf{p}}_{b_{j}}\in\mathbb{R}_{-}$ are the upper and lower limits on battery power. This indicates the discharging and the charging mode of the PCM. The solution to the optimization problems in (12) and (13) are used in calculating the dual variable update. The schematic for this aggregation and distribution setup is provided in Fig. 3. All the nodes present in the system send their current iterate value to the dual updater node and the information of the previous iterate of the dual variable $\boldsymbol{\lambda}_{t}$ is sent to all the nodes. Thus, the dual updater, which acts as a common scheduler to all nodes implements the update law given in (11c). The convergence of the proposed algorithm under the assumed objectives and the constraint sets which are polytopes is well established in the literature ([18]; [19]).
@@ 6 Numerical Simulation
The designed model predictive energy management strategy is tested on a single PGM, PCM, and Load scenario to fully study and analyze the effect of the weights $\beta$ and $\gamma$ on the total energy delivered by the PCM, PGM, and its effect on the PCM capacity loss %. The simulation environment used in this setup is based on MATLAB-Simulink on a real-time simulator. The simulation is run on a desktop with the following configurations Digital Storm Intel Core i9-13900K (5.7GHz Turbo) with 64GB RAM acting as a Host PC and a real-time target with Intel Core 3.6GHz 8-core and 32GB RAM (shown in Fig. 3). The optimization problem presented in (8) with $n_{g}=1$ and $n_{b}=1$ is implemented to study the trade-off characteristics. The parameters used are based on the documentation and the notional SPS component sizing provided by the Office of Naval Research (ONRs) ESRDC ([3]). For all the results presented, the prediction horizon is fixed at ($h=5$seconds). The simulation was run at the fixed step of $10^{-3}$seconds. The rate transition or the communication (measurement exchange) delay between the model and the MPC EM layer was set at $1$second to facilitate optimization convergence.